OpenStudy (anonymous):

Approximate the real number solution(s) to the polynomial function f(x) = x3 + 2x2 – 5x – 6 .

OpenStudy (anonymous):

i really dont Know sorry

OpenStudy (anonymous):

im taking a practice test on flvs right now and im so comfused. i read the lessons adn took notes, but nothigs helping. im about to google it grrr

OpenStudy (anonymous):

f(x) factorises to (x+1)(x+3)(x-2), can you solve the equation now?

OpenStudy (anonymous):

i think so. you set each set of parenthesis equal to zero adn solve right?

OpenStudy (anonymous):

correct, so you should get x = -1, -3, 2

OpenStudy (anonymous):

when f(x) = 0 that is

OpenStudy (anonymous):

thank you sooooooo much!!!!!!

hero (hero):

I don't like these kind of questions because it doesn't necessarily say that f(x) = 0. In my opinion, it should be interpreted as: input a value for f(x) such as x =0,1,2 and simplify the right side: f(0) = (0)^3 + 2(0)^2 + 5(0) - 6 f(1) = (1)^3 + 2(1)^2 + 5(1) - 6 f(2) = (2)^3 +2*(2)^2-5(2) - 6

OpenStudy (anonymous):

i know i dont like these questions either. actually im not real big on algebra. but i love geometry. :)

OpenStudy (shayaan_mustafa):

No. Hero is wrong. to find out solution we have to set the function equal to zero and solve for x, then we could find three answers of x. Regards, Electronics engineer.

OpenStudy (anonymous):

thank you. i already did the assignment with his answer though and im not sure if i got it right or not.

hero (hero):

Shayaan, it doesn't matter whether I'm right or wrong. I'm expressing the fact that the question never suggests to set f(x) = 0. It says to find the solutions to f(x) = x3 + 2x2 – 5x – 6 You cannot legitimately argue against what I have posted as a solution to f(x) = x3 + 2x2 – 5x – 6 because by definition, a solution to an equality involving a variable is the variable that makes both sides of the equation equal. With that being the case, I can justify that f(0) = (0)^3 + 2(0)^2 + 5(0) - 6 f(1) = (1)^3 + 2(1)^2 + 5(1) - 6 f(2) = (2)^3 +2*(2)^2-5(2) - 6 f(n) = (n)^3 - 2(n)^2 - 5(n) - 6 and really, you can't make a legitimate argument against it whether you're an electronics engineer, a teacher, math professor, or a farmer. The solution I have written by the properties of equality and definition of a function suggest that my solution is also correct.

OpenStudy (anonymous):

honestly i wouldnt know if you are right or wrong but when my assignment gets graded. ill let you know. i can always talk to my teacher abotu this but shes not online. ill try to get back to you.

hero (hero):

I believe it's pretty immature for anyone to outright suggest that my solution is wrong without justification.

hero (hero):

That also includes your teacher shayna. :) If she says it is wrong, she also has to justify why.

OpenStudy (anonymous):

oh i know. i just havent spoken to her about it yet. she usually tels me how to solve the problem and i would have told you what she said. im pretty sure you were right though.

hero (hero):

Well, I wouldn't say that I'm 100% right, however, I'm definitely not wrong.

OpenStudy (anonymous):

lol

OpenStudy (shayaan_mustafa):

Here I want to say some thing. Hero is still wrong. As Hero said,"Shayaan, it doesn't matter whether I'm right or wrong. I'm expressing the fact that the question never suggests to set f(x) = 0. It says to find the solutions to f(x) = x3 + 2x2 – 5x – 6." Now it is a matter that you are doing wrong or right and also this question suggest whether to keep f(x)=0 or not. Don't you know how? Regards, Electronics Engineer.